OpenStudy (anonymous):

How do you find the right and left hand derivatives of a piecewise function?

OpenStudy (anonymous):

just take them

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2& \text{if} & x <5 \\ 2x+3& \text{if} & x\geq 5 \end{array} \right. \] derivative is \[f'(x) = \left\{\begin{array}{rcc} 2x& \text{if} & x <5 \\ 2& \text{if} & x>5 \end{array} \right. \]

OpenStudy (anonymous):

derivative from the left at \(x=5\) is \(2\times 5=10\) while the derivative from the right is \(2\)